customer borrows 4000 under a consumer credit load,repayments are calculated to give an APR of 15.4%.Instalments are paid monthly in arrears for 5 years.Calculate the flat rate of intrest


1
Expert's answer
2021-05-12T03:50:41-0400

Loan amount P=4000

Interest rate per annum r=15.4%


Total number of instalments "= 5\times 12=60"


Flat rate of interest "=\dfrac{ P\times r\times 5}{60}=\dfrac{4000\times 15.4\times 5}{60}=5133.3"